OpenStudy (anonymous):

Andrea can clean a house 4 times as fast as Denise. When they work together, Andrea and Denise can clean a large house in 8 hours. How many hours would it take Denise to clean the house by herself? 20 15 10 5

OpenStudy (muzzack):

8*4

OpenStudy (muzzack):

8*4/2

OpenStudy (anonymous):

16?

OpenStudy (muzzack):

B

OpenStudy (anonymous):

15? Howd you get that

OpenStudy (muzzack):

hmm i really domnt know if Andrea can clean a house 4 times as fast as Denise. right and they both do it 8 hours then e.e

OpenStudy (anonymous):

yah that's why its so confusing :P

OpenStudy (muzzack):

well lets say that denise did like 2 hours and andrea did 4 times

OpenStudy (anonymous):

yah

OpenStudy (muzzack):

I think its C

OpenStudy (anonymous):

ok @IMStuck do you have anything or no

OpenStudy (imstuck):

I'm working on it right now. Give me just a minute or two. Please.

OpenStudy (anonymous):

ok great thanks :)

OpenStudy (anonymous):

20

OpenStudy (anonymous):

how

OpenStudy (anonymous):

\[8*\frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

oh ok I guess

OpenStudy (anonymous):

1+4=5/2 x8

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

is it helped

OpenStudy (imstuck):

GOT IT!!!!!

OpenStudy (anonymous):

what is it?

OpenStudy (imstuck):

Ok, if Andrea cleans in 1/4 of the time, and it takes denise, say, one hour to clean a house, it would take Andrea 1/4 of an hour, right? So Denise could get 1/8 of the house done in 1/x and Andrea could get 1/8 the house done in 1/4 of 1/x. Here's your set up\[\frac{ 1 }{ x }+\frac{ 1 }{ 4 }(\frac{ 1 }{ x })=\frac{ 1 }{ 8 }\]Doing all that math and finding the common denominator of 8x, you get 10 = x. That's the number of hours it takes Denise to clean the house by herself.

OpenStudy (anonymous):

omg thank you so much so 10 hours

OpenStudy (imstuck):

Yes, 10 hours. So sorry it took so long!

OpenStudy (anonymous):

Refer to the attached where Mathematica is used for the final calculation.

OpenStudy (anonymous):

no problem thanks for helping me in the first place